The maker of a leading brand of low-calorie microwavable food estimated the following demand equation for its product using data from 26 supermarkets around the country for the month of April

1
Expert's answer
2021-03-03T11:06:25-0500

. Compute the elasticity for each independent variable. Note: Write down all of your calculations.


When P= 500, C=600, I=5,500, A=10,000, and M=5000, using the regression equation,

     QD= - 5200 – 42(500) + 20(600) + 5.2(5500) + 0.20(10000) + 0.25(5000) = 17,650

     Price Elasticity = (P/Q) (∆Q/∆P)

     from the regression equation, ∆Q/∆P = -42.

     So, Price Elasticity (Ep) = (P/Q) (-42) (500/17650) = -1.19, Likewise,

     Ec = 20(600/17560) = 0.68

      

     EA= (P/Q) (0.20) (10000/17650) = 0.11

     EI = (P/Q) (5.2) (5500/17650) = 1.62

     EM = (P/Q) (0.25) (5000/17650) = 0.07